OpenStudy (anonymous):

The length of a rectangular prism is 5 more than twice the width w. The volume of the prism is 2w^3w^2 + 5w. What's the simplified expression for the height of the prism

OpenStudy (anonymous):

EDIT: The length of a rectangular prism is 5 more than twice the width w. The volume of the prism is 2w^3 + w^2 + 5w. What's the simplified expression for the height of the prism

OpenStudy (anonymous):

is w width in your fexpression?

OpenStudy (anonymous):

what is fexpression?

OpenStudy (anonymous):

oh expression... duh

OpenStudy (anonymous):

2w^3 + w^2 + 5w is the w stand for width?

OpenStudy (anonymous):

sorry im tired... anyways yes the w stands for width

OpenStudy (anonymous):

The length of a rectangular prism is 5 more than twice the width w. The volume of the prism is 2w^3 + w^2 + 5w. What's the simplified expression for the height of the prism V=f(w) 2w^3 + w^2 + 5w v=w^2*l svl l=v/w^2 V=w^2*(2*w+5) 2w^3 + w^2 + 5w=w^2*(2*w+5) (2*w+5)= (2w^3 + w^2 + 5w)/w^2 w=5/4 I think knowing that that w=5/4 find l :) hope that helps

OpenStudy (anonymous):

I think it might be l=v/(5/4)^2

OpenStudy (anonymous):

just replace h with l

OpenStudy (anonymous):

the answer in my book says : w-1... do i still have to solve to get that?

OpenStudy (anonymous):

I must have made a mistake somewhere :(

OpenStudy (anonymous):

hm, well you helped me understand how to try and solve it a little more... thats for the help :) I'll try again to figure it out on my own

OpenStudy (anonymous):

*thanks not thats haha

OpenStudy (anonymous):

ah you know the fault was actully on me.... I got the question wrong... its 2w^3 + 7w^2 + 5w.... i forrgot the 7 -_- *facepalm*

OpenStudy (anonymous):

i think in that case w=-5/2 seems wrong tho since it is a nagative :(

OpenStudy (anonymous):

hmm... i dunno...

OpenStudy (anonymous):

do you know what 2w^3 + 7w^2 + 5w would be simplified all by itself?

OpenStudy (anonymous):

w (2 w + 5) (w + 1)

OpenStudy (anonymous):

factor out

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

the answer shoule be 2 w + 5 or 5 +1

OpenStudy (anonymous):

yeah... i dont know how my book got there answer but...

OpenStudy (anonymous):

IDK maybe I need rest and try it in morning :P

OpenStudy (anonymous):

haha yeah me up at an ungodly hour trying to do algebra is not a good mix